one number is 12 more than another the sum of the smaller number and twice the larger number is 39 find the larger number let x be the larger number and y be the smaller number
Answer by SandySharma(35)

let x be the larger number and y be the smaller number.
one number is 12 more than another
x = y + 12
The sum of the smaller number and twice the larger number is 39.
y + 2x = 39

Substitute x in this equation
y + 2 (y +12) = 39
y + 2y + 24 = 39
3y + 24 =39
Subtract 24 on both sides, you will get
3y = 39 -24
3y =15
Divide by 3, you will get
y = 5
x = y + 12 = 5 + 12 = 17
Smaller Number = 5
Larger Number = 17